The reading speed of second grade students is approximately normal, with a mean of 90 words per minute (wpm) and a standard dHaving some trouble and e and f
Answer #1

e)

for normal distribution z score =(X-μ)/σ
here mean=       μ= 90
std deviation   =σ= 10
sample size       =n= 20
std error=σ=σ/√n= 2.2361
probability = P(X>92.2) = P(Z>0.98)= 1-P(Z<0.98)= 1-0.8365= 0.1635

option B: .......92.2 wpm or more is 0.1635 . ......not abundantly more effective,,,,,,,,

f)

for 95th percentile critical value of z= 1.645
therefore corresponding value=mean+z*std deviation= 93.7